OpenStudy (babynini):

f(x)= x/(1+x) g(x) = sin2x Find the following (along with their domains) a) fog b) gof c)fof d) gog

OpenStudy (babynini):

@jim_thompson5910 help please!

OpenStudy (babynini):

a) fog f(g(x)) = f(sin2x)=sin2(1/(1+x))

zepdrix (zepdrix):

Hmm that fog looks a little mixed up :d

zepdrix (zepdrix):

\[\large\rm f(\color{orangered}{x})=\frac{\color{orangered}{x}}{1+\color{orangered}{x}}\]Becomes:\[\large\rm f(\color{orangered}{g(x)})=\frac{\color{orangered}{\sin2x}}{1+\color{orangered}{\sin2x}}\]Ya? :o

zepdrix (zepdrix):

What you posted: \(\large\rm sin\left(2\color{orangered}{\frac{1}{1+x}}\right)\) This is actually the function f being plugged into g, \[\large\rm \sin\left(2\color{orangered}{\frac{1}{1+x}}\right)=\sin\left(2\color{orangered}{f(x)}\right)=g(\color{orangered}{f(x)})\]Where,\[\large\rm \sin(2\color{orangered}{x})=g(\color{orangered}{x})\]

OpenStudy (babynini):

Just a second, trying to work it out on paper :)

OpenStudy (babynini):

haha yeah I wrote it correctly on paper but did it wrong on here xD sorry! just a moment.

zepdrix (zepdrix):

:3

zepdrix (zepdrix):

For part A, do you understand how to find the domain of the composition?

OpenStudy (babynini):

I am only having issues with the domain for c now. The rest are all entered and correct.

OpenStudy (babynini):

c) x/1+2x Domain: (-infinity, -1/2) union (-1/2, infinity) it keeps saying the domain is wrong?

zepdrix (zepdrix):

Oh sorry I ran off for a sec >.<

OpenStudy (babynini):

You're all good :)

zepdrix (zepdrix):

Hmm, I forget how these compositions work sometimes. I guess we need to carry this information around with us before simplifying,\[\large\rm f(x)=\frac{x}{1+x},\qquad\qquad x\ne-1\]So when we get to this composition we have,\[\large\rm f(f(x))=\frac{x}{1+2x},\qquad\qquad x\ne-1,-\frac{1}{2}\]Try that maybe? :o

OpenStudy (babynini):

aaaah. so in interval notation it would be: (-infinity, -1) union (-1,-1/2)union(-1/2, infinity)

zepdrix (zepdrix):

Mmmmm yah I think so :)

zepdrix (zepdrix):